OpenStudy (anonymous):

2500 = x(0.045) + y(0.06) find x, find y

OpenStudy (e.mccormick):

Is this supposed to be in parametric form, or is there a second equation?

OpenStudy (anonymous):

guess so? its just 2500 = x(0.045) + y(0.06)

OpenStudy (e.mccormick):

Hmmm.... because all that does is define a singe line. Perhaps they want it solved for x and solved for y.

OpenStudy (anonymous):

yup

OpenStudy (e.mccormick):

OK. Then you need to basically work with your order of operations, but from bottom up!

OpenStudy (anonymous):

uh...?

OpenStudy (e.mccormick):

Well, you know the order of operations? PEMDAS?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so 2500 = x(.045) + y(.06) x = -(2500-0.6y)/0.045 right?

OpenStudy (e.mccormick):

hmmmm.... \(2500 = x(.045) + y(.06)\implies \\ 2500-y(.06) = x(.045) + y(.06)-y(.06)\implies \\ 2500-y(.06) = x(.045)\implies \\ \left[2500-y(.06)\right]/(.045) = x(.045)/(.045)\implies \\ \left[2500-y(.06)\right]/(.045) = x \) \(\therefore x=\cfrac{2500-y(.06)}{.045}\)

OpenStudy (e.mccormick):

Not sure how you got the - in front.

OpenStudy (anonymous):

I tried that but it x turned to zero(0) (and so did y)

OpenStudy (e.mccormick):

Well, an x of 0 does exist in that solution. So does a y of 0. That is, in that form, the answer for x.

OpenStudy (anonymous):

but it wouldn't be equal to 2500

OpenStudy (e.mccormick):

2500+0=2500 In this particular case, because of the decimals, the intercepts are at much larger numbers. This is the equation of a line. http://www.wolframalpha.com/input/?i=2500+%3D+x%280.045%29+%2B+y%280.06%29&dataset=

OpenStudy (anonymous):

oh, thank you :)

OpenStudy (e.mccormick):

So to find y, just do the same sort of moving things around. If there was a second equation for a second line, and they crossed, there would be one x and one y. But with just one equation, all you have is a line.

OpenStudy (anonymous):

can x be also be found using derivatives?

OpenStudy (e.mccormick):

All you will get with a derivative is the slope.

OpenStudy (anonymous):

oh, thank you:)